# MetroGS: Efficient and Stable Reconstruction of Geometrically Accurate High-Fidelity Large-Scale Scenes

### v1

@@ Abstract
Figure 1: Illustration of the superiority of our method. (a) Our method accurately reconstructs the geometric structure of large-scale urban scenes, faithfully restoring fine details such as buildings, vegetation, and roads. (b) Compared with the SOTA method CityGSV2 [liu2024citygaussianv2], our result are more complete and geometrically precise. (c) Benefiting from a well-designed training framework, our method achieves superior convergence speed and geometric quality. On four RTX 3090 GPUs, our method reaches better performance with less than 25% of the training time required by CityGSV2. The quantitative results are reported based on the quality metrics of Modern Building [xiong2024gauu].
@@ 1 Introduction
3D scene reconstruction is an essential topic in computer vision and graphics. Achieving large-scale, high-precision 3D modeling serves as the foundational support for numerous applications, such as aerial surveying [delplanque2024will, gu2023ue4], autonomous driving [geiger2012we, caesar2020nuscenes, li2024gradiseg], immersive AR/VR [wu20244d, chen2025dual, chen2025learning]. Recent achievements in 3D Gaussian Splatting (3DGS) have notably accelerated the translation of this goal toward practical application [kerbl20233d, yu2024mip, charatan2024pixelsplat, chen2024mixedgaussianavatar], demonstrating remarkable rendering efficiency and visual fidelity. However, current methods, while excelling in rendering quality [li2025stdr, yuan2025robust, liu2025holistic, lin2024vastgaussian], improvements in geometric reconstruction remain relatively limited, leading to an imbalance between visual fidelity and geometric accuracy. This imbalance highlights the need for a scalable reconstruction framework capable of preserving geometric accuracy under large-scale conditions.
@@
In real-world urban environments, large-scale 3D reconstruction must cope with multiple challenges, including objects with diverse structures and scales, varying illumination conditions, and other complex factors. Most existing methods primarily focus on scaling up 2DGS [huang20242d] or PGSR [chen2024pgsr] frameworks, yet their geometric optimization strategies remain underdeveloped. Some methods [liu2024citygaussianv2, lin2025longsplat] rely solely on single-view constraints, making it difficult to maintain structural consistency. While others [chen2024gigags, li2025ulsr, chen20253d] adopt multi-view consistency constraints but typically use single-scale photometric constraints or simple reprojection errors, resulting in limited adaptability to complex large-scale environments. Moreover, illumination and exposure inconsistencies are common in large-scale datasets [zhang2025ref, martin2021nerf], forcing models to reconcile appearance variations during optimization, which compromises geometric consistency. Conventional multi-view consistency optimization struggles to effectively address such issues. Meanwhile, we observe that insufficient initial sampling in weakly textured or sparsely observed regions is another key factor affecting geometric quality, often leading to inaccurate recovery of local structures and resulting in surface holes or structural artifacts.
@@
To overcome these challenges, we propose MetroGS, a novel Gaussian Splatting framework that focuses on achieving Multi-view Efficient Tuning for Robust Optimization in complex urban environments. Specifically, we adopt 2DGS as the core representation for modeling 3D geometry and utilize a distributed training strategy [zhao2024scaling] to efficiently support large-scale scene reconstruction. Building upon this foundation, we introduce a structured dense enhancement scheme. During initialization, the training images are partitioned based on SfM-derived priors, and a pre-trained pointmap model [wang2025pi] is employed to perform dense enhancement on the initial point cloud. In the subsequent densification stage, an additional sparse-compensation mechanism is incorporated to recover the remaining incomplete regions, thereby improving the overall reconstruction completeness and quality. Furthermore, we propose a progressive hybrid geometric refinement strategy. During the early stage of training, we perform a lightweight monocular geometric optimization guided by priors from an off-the-shelf depth estimator [wang2025moge]. As training progresses, a multi-view refinement is introduced in which a carefully designed PatchMatch-based method refines the rendered depths, and the refined maps are further completed with monocular priors to obtain accurate and complete depth maps for subsequent fine-grained geometric optimization. This progressive design effectively balances geometric accuracy and computational efficiency. Finally, we address appearance inconsistencies by decoupling geometry and appearance. Specifically, we introduce a depth-guided appearance modeling module that adopts a Tri-Mip [hu2023tri] structure to store spatial features of the scene. By leveraging the high-quality optimized depth results, the module queries geometry-aligned 3D-consistent feature representations, thereby achieving efficient and stable appearance decoupling. Overall, these components form an efficient and consistent framework for large-scale scene reconstruction, and extensive experiments on multiple large-scale datasets validate its effectiveness.
@@ 2.1 Novel View Synthesis
Novel view synthesis aims to generate high-fidelity images from arbitrary viewpoints by learning an underlying three-dimensional scene representation. The pivotal work NeRF [mildenhall2021nerf] implicitly models the scene using MLPs to encode color and density information for 3D points and viewing directions, enabling novel view synthesis of complex scenes. To address performance limitations, methods exemplified by Tri-MipRF [hu2023tri] employed more advanced feature encoding techniques [muller2022instant, chen2022tensorf] to improve both efficiency and rendering quality. More recently, 3DGS [kerbl20233d] emerged as another influential framework, which models scenes using explicit 3D Gaussian primitives and achieves real-time rendering through differentiable rasterization. Following this breakthrough, a series of works [zhang2024pixel, fang2024mini, ye2024absgs] primarily focus on enhancing rendering quality. These advances collectively inspire the design of our proposed algorithm.
@@ 2.2 Surface Reconstruction
The ability of surface reconstruction to generate accurate 3D geometry from diverse inputs is critical for realizing practical 3D technology. Recently, many advanced methods [guedon2024sugar, yu2024gaussian, wang2024gaussurf] have been developed extended from 3DGS. PGSR [chen2024pgsr] achieves high-fidelity and efficient surface reconstruction by introducing a planar Gaussian representation combined with unbiased depth rendering and multi-view geometric regularization. 2DGS [huang20242d] enhances geometric accuracy by substituting 3D Gaussians with surface-oriented 2D surfels, addressing multi-view inconsistency inherent in 3DGS, and serves as the foundational approach adopted by the best current surface reconstruction methods [huang2025fatesgs, wu2025sparse2dgs, zhang2025ref]. Nevertheless, these approaches are mainly optimized for object-level scenes and cannot be directly applied to large-scale scenes with reliable performance.
@@ 2.3 Large Scale Scene Reconstruction
The task of large-scale reconstruction demands coping with vast amounts of data and more complex scene environments. Several recent works [ren2024octree, lin2024vastgaussian, zhao2024scaling, liu2025holistic, yuan2025robust] have extended 3DGS to large-scale scenes, focusing on rendering quality and efficiency improvements. In contrast, the exploration dedicated to surface reconstruction remains at a relatively early stage. CityGSV2 [liu2024citygaussianv2] continued the strategy of partitioned parallel training, optimizing 2DGS to adapt it for large-scale scene reconstruction, and simultaneously established standard geometric benchmarks for large-scale scenes. CityGS-$\mathcal{X}$ [gao2025citygs] introduced a scalable architecture supporting multi-GPU parallel rendering, and jointly optimizes the scene’s geometry and appearance through batch-level multi-task training. While other methods [gao2024cosurfgs, chen2024gigags] have extended surface reconstruction algorithms to large-scale scenes, their simple geometric optimization struggles with stability in complex large scenes. Our method advances this field by introducing targeted geometric optimization for more robust and higher-quality outcomes.
@@ 3 Preliminaries
3D Gaussian Splatting [kerbl20233d] models a scene as anisotropic Gaussian primitives, each defined by its center, covariance, opacity, and SH coefficients for view-dependent color. Rendering is performed via front-to-back $\alpha$-blending of the $\alpha$-weighted contributions along each ray:
@@
2D Gaussian Splatting [huang20242d] flattens the 3D ellipsoid volume into 2D planar disks, making the primitives highly suitable for explicit 3D surface representation and optimization. In 2DGS, depth calculation is primarily divided into mean depth and median depth. The latter is considered more robust, as it utilizes visibility and treats $T_{i}=0.5$ as the pivot for surface and free space:
@@ 4.1 Scalable Parallel Strategy
We extend 2DGS into a Gaussian-wise distributed training paradigm inspired by the parallel concepts in [zhao2024scaling]. Specifically, the initialization point cloud is uniformly distributed across multiple GPUs for local Gaussian initialization, and multi-view batched training is employed to evenly assign images among devices. Each worker leverages the spatial locality of Gaussian Splatting to fetch only the required Gaussian subsets, enabling efficient communication. During dynamic densification, load balance is maintained through periodic Gaussian redistribution. This distributed design maximizes computational resource utilization and demonstrates excellent scalability, allowing efficient support for large-scale scene reconstruction.
@@ 4.2 Structured Dense Enhancement
Gaussian initialization is based on 3D points from SfM [schonberger2016structure]. However, even in large-scale scenes with dense coverage, the presence of sparse-view or weak-texture regions leads to an overly sparse initial point cloud. To mitigate this, we introduce a structured dense enhancement scheme that separately optimizes initialization and densification.
@@ 4.2.1 Pointmap Model Assisted Initialization
We incorporate the pointmap model [wang2025pi] to obtain auxiliary initial dense point clouds for Gaussian initialization, leveraging its capability for efficient 3D structure prediction. We first construct an undirected image graph $G=(V,E)$, where each node represents an image and each edge weight $w_{ij}$ corresponds to the number of inter-image feature matches estimated by SfM. The graph is partitioned into $N$ clusters, matching the number of available GPUs, by minimizing the normalized cut objective:
@@ 4.3.1 Single-View Optimization
Following [liu2024citygaussianv2], we employ a pretrained depth estimation model [wang2025moge] to obtain a monocular depth prior. The estimated inverse depth is first aligned with the sparse SfM depth, and the L1 loss between the rendered and estimated inverse depths is formulated as $\mathcal{L}_{d}$ to guide depth supervision. In addition, we preserve the depth–normal consistency loss $\mathcal{L}_{n}$ from 2DGS [huang20242d] to further enhance geometric fidelity. In practice, we also observe that large-scale Gaussians often introduce noticeable visual artifacts and blur local details, and their extensive coverage on the image plane leads to heavy GPU memory consumption during training and slows down the optimization process. To mitigate these issues, we introduce a scale regularization term defined as:
$\mathcal{L}_{s}=\frac{1}{|M|}\sum_{i\in M}max(maxs(s_{i})-\tau_{s},\epsilon),$ (6)
@@ 4.3.2 Hybird Multi-View Refinement
Table 1: Comparison with SOTA reconstruction methods on the GauU-Scene [xiong2024gauu] dataset. P and R indicate the Precision and Recall with respect to the ground-truth point cloud. Results highlighted in red, orange, and yellow correspond to the best, second-best, and third-best performances, respectively. “NaN” means no results due to NaN error. “FAIL” means fail to extract meaningful mesh.
@@
Figure 4: Qualitative comparison on the MatrixCity [li2023matrixcity] dataset. Image rendering and mesh reconstruction are compared between our method and CityGSV2 [liu2024citygaussianv2].
@@ 4.4 Depth-Guided Appearance Modeling
Previous works [lin2024vastgaussian, zhang2025ref] show that accurate appearance modeling is crucial for realistic reconstruction, since geometry-only methods often struggle under complex imaging conditions. Existing appearance methods typically do not leverage geometric information, while our method provides high-quality rendered depth that offer a reliable structural prior for appearance learning. Based on this, we design a depth-guided appearance modeling module to ensure appearance estimation under precise geometric constraints, enabling true geometry–appearance decoupling.
Specifically, we employ a Tri-Mip [hu2023tri] structure to store scale-adaptive, multi-resolution 3D features of the scene, which maintain cross-view consistency in space. Given the rendered depth map $D_{r}$, we query the Tri-Mip feature planes using the 3D coordinates of each pixel’s projection, resulting in structure-aligned representations $f_{\text{Tri}}(x)$ . These features provide a stable geometric foundation for appearance estimation, enabling it to focus on color and lighting variations that are independent of geometry. Additionally, each training image $\mathcal{I}_{i}$ is assigned a learnable appearance embedding $l_{i}\in\mathbb{R}^{d}$ to capture global illumination and exposure conditions. The queried Tri-Mip feature $f_{\text{Tri}}(x)$ and the embedding $l_{i}$ are concatenated and passed through a lightweight MLP tone mapper $\mathcal{F}_{\theta}$:
@@
Table 2: Comparison on the MatrixCity  [li2023matrixcity] dataset. “–” indicates the metric was not reported in the original paper.
@@ 4.5 Training Loss
Figure 5: Qualitative results on the GauU-Scene [xiong2024gauu] dataset. We present the image and depth rendering results of our method compared with state-of-the-art methods.
@@ 5.1 Experimental Setup
We conduct comprehensive experiments on the GauU-Scene [xiong2024gauu] dataset and the synthetic MatrixCity [li2023matrixcity] dataset. All experiments are performed on a workstation equipped with four RTX 3090 GPUs. Both datasets provide high-precision ground-truth point clouds, making them reliable benchmarks for evaluating geometric quality in large-scale scene reconstruction. Following the settings in [liu2024citygaussianv2], we evaluate our method on both the aerial-view and street-view versions. The aerial-view images are downsampled to a 1600-pixel long side, while the street-view images retain their original $1000\times 1000$ resolution. For mesh extraction, we adopt the 2DGS methodology, integrating median depth with TSDF fusion. Detailed training and evaluation configurations are provided in the supplementary materials.

### v4

@@ Abstract
Figure 1: Illustration of the superiority of our method. (a) Our method accurately reconstructs the geometric structure of large-scale urban scenes, faithfully restoring fine details such as buildings, vegetation, and roads. (b) Compared with the SOTA method CityGSV2 [34], our result are more complete and geometrically precise. (c) Benefiting from a well-designed training framework, our method achieves superior convergence speed and geometric quality. On four RTX 3090 GPUs, our method reaches better performance with less than 25% of the training time required by CityGSV2. The quantitative results are reported based on the quality metrics of Modern Building [48].
@@ 1 Introduction
3D scene reconstruction is an essential topic in computer vision and graphics. Achieving large-scale, high-precision 3D modeling serves as the foundational support for numerous applications, such as aerial surveying [11, 17], autonomous driving [16, 2, 25], immersive AR/VR [46, 7, 8]. Recent achievements in 3D Gaussian Splatting (3DGS) have notably accelerated the translation of this goal toward practical application [23, 51, 3, 9], demonstrating remarkable rendering efficiency and visual fidelity. However, current methods, while excelling in rendering quality [26, 53, 32, 30], improvements in geometric reconstruction remain relatively limited, leading to an imbalance between visual fidelity and geometric accuracy. This imbalance highlights the need for a scalable reconstruction framework capable of preserving geometric accuracy under large-scale conditions.
@@
In real-world urban environments, large-scale 3D reconstruction must cope with multiple challenges, including objects with diverse structures and scales, varying illumination conditions, and other complex factors. Most existing methods primarily focus on scaling up 2DGS [20] or PGSR [5] frameworks, yet their geometric optimization strategies remain underdeveloped. Some methods [34, 29] rely solely on single-view constraints, making it difficult to maintain structural consistency. While others [6, 28, 10] adopt multi-view consistency constraints but typically use single-scale photometric constraints or simple reprojection errors, resulting in limited adaptability to complex large-scale environments. Moreover, illumination and exposure inconsistencies are common in large-scale datasets [55, 36], forcing models to reconcile appearance variations during optimization, which compromises geometric consistency. Conventional multi-view consistency optimization struggles to effectively address such issues. Meanwhile, we observe that insufficient initial sampling in weakly textured or sparsely observed regions is another key factor affecting geometric quality, often leading to inaccurate recovery of local structures and resulting in surface holes or structural artifacts.
@@
To overcome these challenges, we propose MetroGS, a novel Gaussian Splatting framework that focuses on achieving Multi-view Efficient Tuning for Robust Optimization in complex urban environments. Specifically, we adopt 2DGS as the core representation for modeling 3D geometry and utilize a distributed training strategy [58] to efficiently support large-scale scene reconstruction. Building upon this foundation, we introduce a structured dense enhancement scheme. During initialization, the training images are partitioned based on SfM-derived priors, and a pre-trained pointmap model [45] is employed to perform dense enhancement on the initial point cloud. In the subsequent densification stage, an additional sparse-compensation mechanism is incorporated to recover the remaining incomplete regions, thereby improving the overall reconstruction completeness and quality. Furthermore, we propose a progressive hybrid geometric refinement strategy. During the early stage of training, we perform a lightweight monocular geometric optimization guided by priors from an off-the-shelf depth estimator [44]. As training progresses, a multi-view refinement is introduced. Inspired by [1, 49], we adopt a carefully designed PatchMatch-based method to refine the rendered depths, and further complete the refined depth maps with monocular priors to obtain accurate and complete depth estimates for subsequent fine-grained geometric optimization. This progressive design effectively balances geometric accuracy and computational efficiency. Finally, we address appearance inconsistencies by decoupling geometry and appearance. Specifically, we introduce a depth-guided appearance modeling module that adopts a Tri-Mip [19] structure to store spatial features of the scene. By leveraging the high-quality optimized depth results, the module queries geometry-aligned 3D-consistent feature representations, thereby achieving efficient and stable appearance decoupling. Overall, these components form an efficient and consistent framework for large-scale scene reconstruction, and extensive experiments on multiple large-scale datasets validate its effectiveness.
@@ 2.1 Novel View Synthesis
Novel view synthesis aims to generate high-fidelity images from arbitrary viewpoints by learning an underlying three-dimensional scene representation. The pivotal work NeRF [37] implicitly models the scene using MLPs to encode color and density information for 3D points and viewing directions, enabling novel view synthesis of complex scenes. To address performance limitations, methods exemplified by Tri-MipRF [19] employed more advanced feature encoding techniques [38, 4] to improve both efficiency and rendering quality. More recently, 3DGS [23] emerged as another influential framework, which models scenes using explicit 3D Gaussian primitives and achieves real-time rendering through differentiable rasterization. Following this breakthrough, a series of works [56, 13, 50] primarily focus on enhancing rendering quality. These advances collectively inspire the design of our proposed algorithm.
@@ 2.2 Surface Reconstruction
The ability of surface reconstruction to generate accurate 3D geometry from diverse inputs is critical for realizing practical 3D technology. Recently, many advanced methods [18, 52, 42] have been developed extended from 3DGS. PGSR [5] achieves high-fidelity and efficient surface reconstruction by introducing a planar Gaussian representation combined with unbiased depth rendering and multi-view geometric regularization. 2DGS [20] enhances geometric accuracy by substituting 3D Gaussians with surface-oriented 2D surfels, addressing multi-view inconsistency inherent in 3DGS, and serves as the foundational approach adopted by the best current surface reconstruction methods [21, 47, 55]. Nevertheless, these approaches are mainly optimized for object-level scenes and cannot be directly applied to large-scale scenes with reliable performance.
@@ 2.3 Large Scale Scene Reconstruction
The task of large-scale reconstruction demands coping with vast amounts of data and more complex scene environments. Several recent works [39, 30, 58, 32, 53] have extended 3DGS to large-scale scenes, focusing on rendering quality and efficiency improvements. In contrast, the exploration dedicated to surface reconstruction remains at a relatively early stage. CityGSV2 [34] continued the strategy of partitioned parallel training, optimizing 2DGS to adapt it for large-scale scene reconstruction, and simultaneously established standard geometric benchmarks for large-scale scenes. CityGS-$\mathcal{X}$ [15] introduced a scalable architecture supporting multi-GPU parallel rendering, and jointly optimizes the scene’s geometry and appearance through batch-level multi-task training. While other methods [14, 6] have extended surface reconstruction algorithms to large-scale scenes, their simple geometric optimization struggles with stability in complex large scenes. Our method advances this field by introducing targeted geometric optimization for more robust and higher-quality outcomes.
@@ 3 Preliminaries
3D Gaussian Splatting [23] models a scene as anisotropic Gaussian primitives, each defined by its center, covariance, opacity, and SH coefficients for view-dependent color. Rendering is performed via front-to-back $\alpha$-blending of the $\alpha$-weighted contributions along each ray:
@@
2D Gaussian Splatting [20] flattens the 3D ellipsoid volume into 2D planar disks, making the primitives highly suitable for explicit 3D surface representation and optimization. In 2DGS, depth calculation is primarily divided into mean depth and median depth. The latter is considered more robust, as it utilizes visibility and treats $T_{i}=0.5$ as the pivot for surface and free space:
@@ 4.1 Scalable Parallel Strategy
We extend 2DGS into a Gaussian-wise distributed training paradigm inspired by the parallel concepts in [58]. Specifically, the initialization point cloud is uniformly distributed across multiple GPUs for local Gaussian initialization, and multi-view batched training is employed to evenly assign images among devices. Each worker leverages the spatial locality of Gaussian Splatting to fetch only the required Gaussian subsets, enabling efficient communication. During dynamic densification, load balance is maintained through periodic Gaussian redistribution. This distributed design maximizes computational resource utilization and demonstrates excellent scalability, allowing efficient support for large-scale scene reconstruction.
@@ 4.2 Structured Dense Enhancement
Gaussian initialization is based on 3D points from SfM [40]. However, even in large-scale scenes with dense coverage, the presence of sparse-view or weak-texture regions leads to an overly sparse initial point cloud. To mitigate this, we introduce a structured dense enhancement scheme that separately optimizes initialization and densification.
@@ 4.2.1 Pointmap Model Assisted Initialization
We incorporate the pointmap model [45] to obtain auxiliary initial dense point clouds for Gaussian initialization, leveraging its capability for efficient 3D structure prediction. We first construct an undirected image graph $G=(V,E)$, where each node represents an image and each edge weight $w_{ij}$ corresponds to the number of inter-image feature matches estimated by SfM. The graph is partitioned into $N$ clusters, matching the number of available GPUs, by minimizing the normalized cut objective:
@@ 4.3.1 Single-View Optimization
Following [34], we employ a pretrained depth estimation model [44] to obtain a monocular depth prior. The estimated inverse depth is first aligned with the sparse SfM depth, and the L1 loss between the rendered and estimated inverse depths is formulated as $\mathcal{L}_{d}$ to guide depth supervision. In addition, we preserve the depth–normal consistency loss $\mathcal{L}_{n}$ from 2DGS [20] to further enhance geometric fidelity. In practice, we also observe that large-scale Gaussians often introduce noticeable visual artifacts and blur local details, and their extensive coverage on the image plane leads to heavy GPU memory consumption during training and slows down the optimization process. To mitigate these issues, we introduce a scale regularization term defined as:
$\mathcal{L}_{s}=\frac{1}{|M|}\sum_{i\in M}max(max(s_{i})-\tau_{s},\epsilon),$ (6)
@@ 4.3.2 Hybird Multi-View Refinement
Table 1: Comparison with SOTA reconstruction methods on the GauU-Scene [48] dataset. P and R indicate the Precision and Recall with respect to the ground-truth point cloud. Results highlighted in red, orange, and yellow correspond to the best, second-best, and third-best performances, respectively. “NaN” means no results due to NaN error. “FAIL” means fail to extract meaningful mesh.
@@
Figure 4: Qualitative comparison on the MatrixCity [24] dataset. Image rendering and mesh reconstruction are compared between our method and CityGSV2 [34].
@@ 4.4 Depth-Guided Appearance Modeling
Previous works [30, 55] show that accurate appearance modeling is crucial for realistic reconstruction, since geometry-only methods often struggle under complex imaging conditions. Existing appearance methods typically do not leverage geometric information, while our method provides high-quality rendered depth that offer a reliable structural prior for appearance learning. Based on this, we design a depth-guided appearance modeling module to ensure appearance estimation under precise geometric constraints, enabling true geometry–appearance decoupling.
Specifically, we employ a Tri-Mip [19] structure to store scale-adaptive, multi-resolution 3D features of the scene, which maintain cross-view consistency in space. Given the rendered depth map $D_{r}$, we query the Tri-Mip feature planes using the 3D coordinates of each pixel’s projection, resulting in structure-aligned representations $f_{\text{Tri}}(x)$ . These features provide a stable geometric foundation for appearance estimation, enabling it to focus on color and lighting variations that are independent of geometry. Additionally, each training image $\mathcal{I}_{i}$ is assigned a learnable appearance embedding $l_{i}\in\mathbb{R}^{d}$ to capture global illumination and exposure conditions. The queried Tri-Mip feature $f_{\text{Tri}}(x)$ and the embedding $l_{i}$ are concatenated and passed through a lightweight MLP tone mapper $\mathcal{F}_{\theta}$:
@@
Table 2: Comparison on the MatrixCity  [24] dataset. “–” indicates the metric was not reported in the original paper.
@@ 4.5 Training Loss
Figure 5: Qualitative results on the GauU-Scene [48] dataset. We present the image and depth rendering results of our method compared with state-of-the-art methods.
@@ 5.1 Experimental Setup
We conduct comprehensive experiments on the GauU-Scene [48] dataset and the synthetic MatrixCity [24] dataset. All experiments are performed on a workstation equipped with four RTX 3090 GPUs. Both datasets provide high-precision ground-truth point clouds, making them reliable benchmarks for evaluating geometric quality in large-scale scene reconstruction. Following the settings in [34], we evaluate our method on both the aerial-view and street-view versions. The aerial-view images are downsampled to a 1600-pixel long side, while the street-view images retain their original $1000\times 1000$ resolution. For mesh extraction, we adopt the 2DGS methodology, integrating median depth with TSDF fusion. Detailed training and evaluation configurations are provided in the supplementary materials.